OpenStudy (anonymous):

Find the volume of a square pyramid with base edges of 48 cm and a slant height of 26 cm. A. 11,520 cm3 B. 23,040 cm3 C. 7,680 cm3 D. 768 cm3

OpenStudy (anonymous):

1/3(48)^2(26)=416.......?

OpenStudy (anonymous):

V=1/3(l^2)(h)

OpenStudy (anonymous):

@Phydeaux would u happen to know this , i have the form but dnt know y its not comming out right

OpenStudy (anonymous):

l=48 h=26

OpenStudy (anonymous):

Sooo V= ((48^2)(26))/3

OpenStudy (anonymous):

Hmm. I get 19,968. Give me a sec.

OpenStudy (anonymous):

wut does pl mean again

OpenStudy (anonymous):

found my mistake.... I was using slant height rather than actual height. One sec.

OpenStudy (anonymous):

I got C. The height is 10, so ((48^2)(10))/3=V

OpenStudy (anonymous):

thank u ya i was using 26

OpenStudy (anonymous):

You have to use pythagorean theorem to calculate the height rather than slant height, and use regular height. You're welcome. :]

OpenStudy (anonymous):

 sorry I meant the height to be 10 in my drawing, but here it is!!!